Find the following for f(x) = e^{−x^{2}} + 1 (if they exist; if they don’t exist, state so). Use this information to graph f. Note f′(x) = −2xe^{−x^{2}}, f′′(x) = 2e^{−x^{2}}[2x^{2} − 1] Is f even, odd, periodic? What is the domain and range of f? [1.5] 1a.) critical numbers: [1.5] 1b.) local maximum(s) occur at x = [1.5] 1c.) local minimum(s) occur at x = [1.5] 1d.) The global maximum of f on the interval [0, 5] is and occurs at x = [1.5] 1e.) The global minimum of f on the interval [0, 5] is and occurs at x = [1.5] 1f.) Inflection point(s) occur at x = [1.5] 1g.) f increasing on the intervals [1.5] 1h.) f decreasing on the intervals [1.5] 1i.) f is concave up on the intervals [1.5] 1j.) f is concave down on the intervals [1.5] 1k.) Equation(s) of vertical asymptote(s)

[4] 1l.) Equation(s) of horizontal and/or slant asymptote(s) [4.5] 1m.) Graph f